The lenght of a rectangle is 5cm more than 4 times it's width. The perimeter is 124cm find it's dimension?
Find the modulas 1A multi and argument
Z=3+6i
Z=-3-4i
Z=8-3i
z=-6-3i

Found 2 solutions by stanbon, Alan3354:
Answer by stanbon(75887)

The length of a rectangle is 5cm more than 4 times it's width. The perimeter is 124cm find it's dimension?
--------------
Perimeter = 2(width + length)
124 = 2(w + 4w+5)
62 = 5w+5
5w = 57
w = 57/5
w = 11 2/5 cm (width)
length = 4w+5 = 4(57/5) + 5
= (228/5) + (25/5)
= 253/5
= 50 3/5 cm (length)
==========================
Find the modulas 1A multi and argument
-----
Comment: One of those is the length; the other is the angle.
I'll let you decide which is modulas and which as argument.
=====================================
Z=3+6i
length = sqrt(3^2 + 6^2)
angle = arctan(6/3) = 63.43 degrees
--------------------------------------
Z=-3-4i
Note: This point is in the 3rd quadrant where tangents are
positive but angles are between 180 and 270.
---
length = sqrt(3^2 + 4^2)
angle = arctan(-4/-3) = 53.13 degrees which is 180+53.13 = 233.13
----------------------
Z=8-3i
length = sqrt(8^2+3^2)
angle is in QIV; arctan(-3/8)= -20.56 degrees
-----------------------------------------
z=-6-3i
I'll leave this one to you.
================================
Cheers,
Stan H.

Answer by Alan3354(69443)

it's = it is.
The possessive is "its"